the area of the square is 125 square cm.what is the length of the side of the square?please help...thanks in advance
Answer by addingup(3677)

What is a multiplication? A quick addition:
3+3+3+3+3= 15 Instead of adding 3 five times we can:
3*5=15 and we have replaced four operations with one.
The opposite of multiplication is division.
--------------------
And what is an exponent? A quick multiplication:
3*3*3*3*3= 243 Instead of multiplying 3 five times we can:
3^5
The opposite of an exponent is a root
-------------------------------------
The area of the square is S^2 or what is the same S*S. So we have to undo the square. Since it's S^2 we need the square root:
sqrt125= 11.18 is the length of each side.
-----------------------------------
Check: 11.18*11.18= 125 We have the correct answer